OpenStudy (abhisar):

Find the value of x 800cosx = (800sinx+870)0.55

OpenStudy (abhisar):

@ganeshie8 @phi

OpenStudy (abhisar):

yes...

OpenStudy (abhisar):

u mean 55/100

OpenStudy (abhisar):

steps are gone !

OpenStudy (abhisar):

okay u type it here...i'll type there at the same time

OpenStudy (abhisar):

nopes !

OpenStudy (abhisar):

How this ? \(\huge (\frac{ 80000 }{ 55 }-870)cosx\)

OpenStudy (abhisar):

(A.B)-C = (A-C).B ???

OpenStudy (abhisar):

yes i think so :(

OpenStudy (abhisar):

yeah i did that....then i am struck...how to find x then ?

OpenStudy (anonymous):

@ganeshie8